# Forensics Adapter: Adapting CLIP for Generalizable Face Forgery Detection

### v1

@@ Abstract
We describe the Forensics Adapter, an adapter network designed to transform CLIP into an effective and generalizable face forgery detector. Although CLIP is highly versatile, adapting it for face forgery detection is non-trivial as forgery-related knowledge is entangled with a wide range of unrelated knowledge. Existing methods treat CLIP merely as a feature extractor, lacking task-specific adaptation, which limits their effectiveness. To address this, we introduce an adapter to learn face forgery traces – the blending boundaries unique to forged faces, guided by task-specific objectives. Then we enhance the CLIP visual tokens with a dedicated interaction strategy that communicates knowledge across CLIP and the adapter. Since the adapter is alongside CLIP, its versatility is highly retained, naturally ensuring strong generalizability in face forgery detection. With only trainable parameters, our method achieves a significant performance boost, improving by approximately on average across five standard datasets. We believe the proposed method can serve as a baseline for future CLIP-based face forgery detection methods.
@@ 2 Related Works
Face Forgery Detection. Detecting face forgeries is a significant topic in recent years [37, 51, 47, 11, 3]. With the great advancement of deep learning, recent methods are typically designed on Deep Neural Networks (DNNs) to capture the forgery traces, utilizing features ranging from the biological signals (e.g., eye blinking [37], heartbeat rhythm [50], head pose [69], facial action units [1]), frequency signals (e.g., [45, 51, 41]), and auto-learned clues with dedicated training schemes (e.g., [56, 6, 11]). However, these methods usually rely on specific training datasets, which can lead to overfitting on known data distributions and hinder their ability to generalize across different datasets. To enhance detection generalization, many methods have been proposed to create pseudo-fake faces by simulating blending boundaries in real images [35, 55, 31, 47, 3]. While these methods improve the generalization to a certain extent, limited data diversity still restricts detectors from learning generic features.
@@ 3.2 Interaction between Adapter and CLIP
Absorbing CLIP Visual Tokens. The features of CLIP are highly versatile and can greatly support the learning process of the adapter. Therefore, we fuse the features of CLIP into the adapter. Since CLIP and the adapter have different goals, we limit the feature fusion to the shallow layers of the adapter, avoiding interference with the learning of task-specific features. For a standard 24-layer ViT-L/14 model, we fuse features from layers to the layers of the adapter. Since the feature dimensions of CLIP and the adapter are inconsistent, we employ a convolution to align them. We then add these two features as the fused ones.
Enhancing CLIP Visual Tokens. To make CLIP focus more on blending boundary traces, we convey the knowledge of the adapter to CLIP without modifying any parameters. We achieve this by utilizing an attention bias strategy similar to the one proposed in [63]. Note that CLIP tokens comprise visual tokens and token, denoted as . For each layer, we duplicate copies of token to form an independent token as . The entire tokens at a layer can be written as . Then we update using self-attention based on the output of the adapter and original visual tokens in CLIP, which is defined as
@@ 4.3 Analysis
Where Using Task-specific Learning. Theoretically, patch-wise contrastive learning and sample-wise contrastive learning can be used on both CLIP and adapter. Thus, we study where the learning should added. The results, shown in Table 5, highlight the importance of the proposed task-specific learning strategies. By using the learning on the adapter (ADP) side, we can improve the performance by in AUC and AP while reducing EER by on average. By applying the learning solely on the CLIP side, we can achieve a performance increase by around in AUC and AP while lowering EER by on average. When adding the learning strategy to both CLIP and the adapter, our method achieves the best performance.
Using Various LLMs. This part investigates the effect of using various LLMs. Specifically, we validate three LLMs: BLIP ViT-L/14 [34], CLIP ViT-B/16 [52], and CLIP ViT-L/14 [52] (the one used in main experiment). We observe that BLIP is not suitable for the task, which highly degrades the performance compared to CLIP variants. This is possible because BLIP is trained using only 129 million image-text pairs and designed for image-text retrieval and VQA tasks with a bootstrapping mechanism. In contrast, CLIP is trained on a much larger dataset containing 400 million samples, enabling superior zero-shot capabilities in the modality of images. Straightforwardly, using large ViT can obtain more performance gain than the base version, which aligns with our understanding.
@@
Exploration of the Text Modality. In addition to using the visual encoder of CLIP, we also explore the incorporation of text encoder. Following CoOP [71], we design learnable contexts within the prompts and combine the text tokens with visual tokens. The implementation details are provided in the Supplementary. Table 8 presents the results. Surprisingly, the performance after using text modality is notably lower than only using the visual modality. This may be because the forged faces are mainly reliant on visual representations, and the integration of text modality introduces irrelevant noise, which distracts attention from the correct forgery traces.
@@ 5 Conclusion
In this paper, we introduce Forensics Adapter, an adapter network designed to transform CLIP into a generalizable face forgery detector. Different from existing methods, our adapter is specifically tailored to capture the unique blending boundaries characteristic of forged faces using task-specific objectives. We then introduce an interactive strategy that enables CLIP and the adapter to collaboratively focus on face forgery traces. Notably, our solution of adapting CLIP for face forgery detection is novel, addressing a gap that previous methods have not explored. Extensive experiments demonstrate that, with only 5.7M trainable parameters, our adapter yields substantial performance gains, improving the performance by approximately on average across five standard datasets. We believe our method can serve as a baseline for future CLIP-based face forgery detection techniques.

### v2

@@ Abstract
We describe the Forensics Adapter, an adapter network designed to transform CLIP into an effective and generalizable face forgery detector. Although CLIP is highly versatile, adapting it for face forgery detection is non-trivial as forgery-related knowledge is entangled with a wide range of unrelated knowledge. Existing methods treat CLIP merely as a feature extractor, lacking task-specific adaptation, which limits their effectiveness. To address this, we introduce an adapter to learn face forgery traces – the blending boundaries unique to forged faces, guided by task-specific objectives. Then we enhance the CLIP visual tokens with a dedicated interaction strategy that communicates knowledge across CLIP and the adapter. Since the adapter is alongside CLIP, its versatility is highly retained, naturally ensuring strong generalizability in face forgery detection. With only trainable parameters, our method achieves a significant performance boost, improving by approximately on average across five standard datasets. We believe the proposed method can serve as a baseline for future CLIP-based face forgery detection methods. The code is available at https://github.com/OUC-VAS/ForensicsAdapter.
@@ 2 Related Works
Face Forgery Detection. Detecting face forgeries is a significant topic in recent years [39, 53, 49, 11, 3, 62]. With the great advancement of deep learning, recent methods are typically designed on Deep Neural Networks (DNNs) to capture the forgery traces, utilizing features ranging from the biological signals (e.g., eye blinking [39], heartbeat rhythm [52], head pose [72], facial action units [1]), frequency signals (e.g., [47, 53, 43]), and auto-learned clues with dedicated training schemes (e.g., [58, 6, 11]). However, these methods usually rely on specific training datasets, which can lead to overfitting on known data distributions and hinder their ability to generalize across different datasets. To enhance detection generalization, many methods have been proposed to create pseudo-fake faces by simulating blending boundaries in real images [37, 57, 33, 49, 3]. While these methods improve the generalization to a certain extent, limited data diversity still restricts detectors from learning generic features.
@@ 3.2 Interaction between Adapter and CLIP
Absorbing CLIP Visual Tokens. The features of CLIP are highly versatile and can greatly support the learning process of the adapter. Thus we fuse these features into the adapter. Since CLIP and the adapter have different goals, we limit the feature fusion to the shallow layers of the adapter, avoiding interference with the learning of task-specific features. For a standard 24-layer ViT-L/14 model, we fuse features from layers to the layers of the adapter. Since the feature dimensions of CLIP and the adapter are inconsistent, we employ a convolution to align them. We then add these two features as the fused ones.
Enhancing CLIP Visual Tokens. To make CLIP focus more on blending boundary traces, we convey the knowledge of the adapter to CLIP without modifying any parameters. We achieve this by utilizing an attention bias strategy similar to the one proposed in [66]. Note that CLIP tokens comprise visual tokens and token, denoted as . For each layer, we duplicate copies of token to form an independent token as . The entire tokens at a layer can be written as . Then we update using self-attention based on the output of the adapter and original visual tokens in CLIP, as
@@ 4.3 Analysis
Where Using Contrastive Learning. Theoretically, patch-wise contrastive learning and sample-wise contrastive learning can be used on both CLIP and adapter. Thus, we study where the learning should added. The results, shown in Table 5, highlight the importance of the proposed task-specific learning strategies. By using the learning on the adapter (ADP) side, we can improve the performance by in AUC and AP while reducing EER by on average. By applying the learning solely on the CLIP side, we can achieve a performance increase by around in AUC and AP while lowering EER by on average. When adding the learning strategy to both CLIP and the adapter, our method achieves the best performance.
Using Various LLMs. This part investigates the effect of using various LLMs. Specifically, we validate three LLMs: BLIP ViT-L/14 [36], CLIP ViT-B/16 [54], and CLIP ViT-L/14 [54] (the one used in main experiment). The results are shown in Table 6. We observe that BLIP is not suitable for the task, which highly degrades the performance compared to CLIP variants. This is possible because BLIP is trained using only 129 million image-text pairs and designed for image-text retrieval and VQA tasks with a bootstrapping mechanism. In contrast, CLIP is trained on a much larger dataset containing 400 million samples, enabling superior zero-shot capabilities in the modality of images. Straightforwardly, using large ViT can obtain more performance gain than the base version, which aligns with our understanding.
@@
Using Various Fine-Tuning Methods. We also study the effect of using various fine-tuning methods, including LoRA [24] and subset fine-tuning [22]. As shown in Table 7, without task-specific adaptation, other fine-tuning methods yield only moderate performance.
| Method | CDF-v2 | DFDCP |
| LoRA [24] | 0.797 | 0.800 |
| subset fine-tuning [22] | 0.770 | 0.807 |
| Ours | 0.900 | 0.890 |
@@
Exploration of the Text Modality. In addition to using the visual encoder of CLIP, we also explore the incorporation of text encoder. Following CoOP [74], we design learnable contexts within the prompts and combine the text tokens with visual tokens. Table 9 presents the results. Surprisingly, the performance after using text modality is notably lower than only using the visual modality. This may be because the forged faces are mainly reliant on visual representations, and the integration of text modality introduces irrelevant noise, which distracts attention from the correct forgery traces.
@@ 5 Conclusion
In this paper, we introduce Forensics Adapter, an adapter network designed to transform CLIP into a generalizable face forgery detector. Different from existing methods, our adapter is specifically tailored to capture the unique blending boundaries characteristic of forged faces using task-specific objectives. We then introduce an interactive strategy that enables CLIP and the adapter to collaboratively focus on face forgery traces. Notably, our solution of adapting CLIP for face forgery detection is novel, addressing a gap that previous methods have not explored. Extensive experiments demonstrate that, with only trainable parameters, our adapter yields substantial performance gains, improving the performance by approximately on average across five standard datasets. We believe our method can serve as a baseline for future CLIP-based face forgery detection techniques.
Acknowledgement. This work is supported in part by the National Natural Science Foundation of China (No.62402464), Shandong Provincial Natural Science Foundation (No.ZR2024QF035), and China Postdoctoral Science Foundation (No.2021TQ0314; No.2021M703036).